# Developers’ Perspectives on Software Licensing: Current Practices, Challenges, and Tools

### v1

@@ 3.3. Interview design and analysis
To better understand participants’ responses and gain further insights into their views on software licensing and related issues, we conducted semi-structured follow-up interviews over Zoom. At the conclusion of the survey, we provided respondents with an option to leave an email address if they would be willing to participate in a follow-up interview, which 21 participants utilized. We contacted all of these individuals, which led to seven respondents agreeing and participating in follow-up interviews.
Each interview lasted 20 to 30 minutes and was recorded to facilitate transcription and subsequent analysis. The interviews were transcribed using Whisper (Whisper, [n. d.]) and reviewed against the original recordings for accuracy by one researcher each. Three reviewers assigned topic labels to relevant portions of each transcript, and all three reviewers then discussed the labels to reach a consensus.
To preserve interviewee anonymity, we have assigned each one an identifier to attribute the source of quotations (e.g., $RI_{4}$). As with the survey results, we have corrected errors in grammar for readability and have indicated deletions or additions by ellipses or brackets.
@@ 3.4. Participant demographics
Respondents reported a variety of different backgrounds and experiences. Twenty-one countries were represented, with the largest groups of participants residing in the United States (22, 38%) , Germany (6, 10%), Bangladesh (5, 9%), and Colombia (5, 9%). Forty respondents (69%) reported that they worked on proprietary software, 33 (57%) on open-source projects that relied on third-party components, and 14 (24%) on open-source projects with no dependencies. The presence of these different groups bolsters our results by allowing us to capture information on licensing tasks for different use cases. Most of the respondents described their primary role as programmer (32, 55%), with other common roles including researcher (10, 17%), project technical lead (5, 9%), tester (3, 5%), educator (3, 5%), and software architect (2, 3%). Development experience ranged from less than one year to 35 years, with an average of 9.9 years and a median of 7 years. Domains also varied broadly, ranging from contracted government work and healthcare to game development and banking. Almost two-thirds of the respondents had experience with web development (37, 64%). Other common system types developed by the respondents included libraries/frameworks (25, 43%), desktop applications (20, 34%), mobile applications (15, 26%), development tools (14, 24%), middleware (13, 22%), and AI-intensive systems (9, 16%). Although not extensive, the developers in the response pool represent a broad range of (self-reported) experiences and backgrounds, with likely diverse experiences with the software licensing process.
We evaluated participants’ experience with licensing by asking them if they had any training or education in software licensing, and if so, whether that training was formal (e.g., college classes, degree, or certification) or informal (self-learning, employer-provided, etc.). As shown in Table 1, more than half of the respondents indicated that they had some kind of training (35, 60%), although most reported that it was informal training (25, 43%). We note that our goal was to investigate the perspectives of developers broadly: while licensing experts and compliance specialists can be expected to know more about licensing issues and may have different strategies for addressing them than an average developer, everyday developers are on the front lines of license compliance (Wintersgill et al., 2024). Capturing their perspectives grants insight into how compliance is actually done (or not) by the people who encounter the need for it regularly.
@@ 4. Results
In the following, we present the results of our study, which emerged from the survey responses and interviews collected.
4.1. RQ1: Developer Licensing Challenges
Training Count Percentage
No 23 39.66%
Yes, informal 25 43.10%
Yes, formal 4 6.90%
Yes, formal and informal 6 10.34%
Table 1. Respondents’ license training
This section discusses participants’ responses to questions regarding the different software licensing tasks they perform and the challenges they face.
4.1.1. Most common challenges
Participants were presented with a list of software licensing-related tasks and asked to identify how frequently they engaged in each task on a scale of “never,” “occasionally,” “frequently,” or “all the time.” The tasks were derived from our previous experience working with licensing issues and from prior work cataloging licensing bugs (Vendome et al., 2018). As shown in Figure 3, the tasks developers reported most often engaging in (i.e., frequently, or all the time) were identifying the licenses of software components they depend on, understanding the terms of use/service associated with the tools they use to develop software, and verifying that their software complies with all relevant licenses. Developers reported that they assigned a license to their own software slightly less often, though this was the task most reported by the respondents as being performed “all the time” (10). By contrast, most respondents indicated that they never spent time changing their software’s license (42) or acquiring or granting license exceptions (40), and slightly less than half reported that they never spent time identifying key terms from software licenses (28). Notably, developers reported performing all tasks “never” or “occasionally” more than they reported performing them “frequently” or “all the time,” suggesting that they perform licensing tasks relatively infrequently overall.
@@ 4.1.1. Most common challenges
Changing a software project’s license can be difficult (Wintersgill et al., 2024), particularly for open-source projects, but such changes do occur (Vendome et al., 2017a). Few developers, however, reported having experience with changing the license of a software component, and no respondents reported performing this task “all the time.” Similarly, most developers reported that they had no experience in acquiring or granting license exceptions, in which an exception to the terms of a component’s license is granted to a particular licensee (for example, to grant additional permissions or remove restrictions that would otherwise be in place). License exceptions can be difficult to obtain from open source projects (Wintersgill et al., 2024) and may be a process with which developers are unfamiliar; indeed, this category received the most “I don’t know” responses, as shown in Table 2. Almost half of the respondents reported that they never spent time identifying key terms from licenses. This may suggest that developers spend little time analyzing the contents of software licenses, or that they feel they understand available software licenses and how they interact, so they do not need to spend further time reviewing them.
@@
4.1.2. Most difficult challenges
Developers were also asked to rate the difficulty of licensing tasks using a 5-point Likert scale from “extremely difficult” to “extremely easy.” Here, as shown in Figure 4, we observe that tasks such as identifying software licenses, understanding terms associated with development tools, and assigning a license were viewed as easiest, while changing a license, dealing with license exceptions, and identifying key license terms received more mixed responses. Notably, respondents did not reach a consensus for any task, showing that it was easy: the results show mixed opinions at best. Our results confirm previous work showing that developers can find licensing tasks challenging (Almeida et al., 2017, 2019).
Finding 3: Developers gave mixed answers regarding the difficulty of most licensing tasks examined. No tasks were broadly considered easy.
@@ 4.1.2. Most difficult challenges
Verifying whether software complies with all relevant licenses was cited as the most difficult task by respondents, being the task most frequently labeled “somewhat difficult” (25) as well as most frequently labeled “extremely difficult” (13). This is particularly notable, as compliance verification was reported as one of the most frequently performed tasks in the previous question. Given that this task can occur whenever components are added or changed, it is possible that its perceived difficulty dissuades developers from performing it. Identifying key license terms was also frequently cited as somewhat/extremely difficult. That developers struggle to understand licenses and their terms, and that verifying compliance proves a difficult task for them, is a widely accepted truth in this area (Almeida et al., 2019; Gangadharan et al., 2008, 2012; Vendome et al., 2018). Our results confirm these previous findings while providing additional depth by demonstrating which tasks developers find most challenging among these, and contextualizing them with details on how often these challenging issues arise during development. This further motivates the need for research efforts to improve the process of confirming software compliance.
Finding 4: Verifying license compliance, while one of the most common licensing tasks, was also reported to be the most difficult. Identifying key license terms was also frequently perceived as “extremely difficult.”
@@
Table 2. “I don’t know” / “I have no opinion” responses to task frequency and difficulty questions
@@ 4.1.3. Licensing factors
To better understand the elements involved in the decision to license software, participants were asked to identify the factors they or their organizations consider when assigning licenses to software. Participants could select multiple such factors. As shown in Figure 5, developers most commonly cited legal considerations surrounding their software (30) as influencing the license assigned to their software. Together with the second most popular answer, organizational policies (24), this illustrates how licensing decisions are often made for non-technical reasons. The third most frequent answer, dependency licenses (23), shows how constraints based on third-party components can similarly affect such decisions. Only beyond this point do we observe specific considerations related directly to development-focused concerns, such as the intended distribution of the software project (21) or its intended use (19). Philosophical considerations, such as the participant’s or their organization’s views on ethical software distribution and use (13) or their views on the nature or purpose of OSS development (12), received fewer responses. Overall, these results highlight a hierarchy of considerations when choosing a license: first, any legal obligations or organizational policies must be adhered to, then technical limitations based on the project and how it should be used, followed by philosophies regarding software and its reuse. It should be noted that as participants were able to select multiple responses, a given respondent could highlight multiple of these groupings to be considered in tandem with each other, and for some respondents, the categories may overlap. However, the preeminence of non-engineering or non-usage requirements illustrates the first priorities respondents reported having when making licensing decisions. As such policies may vary among jurisdictions or among organizations, we suggest that licensing tools should accommodate organizational choices.
Finding 5: When considering how to license software, the surveyed developers appeared to first consider legal and organizational obligations before applying technical constraints based on the project and then, less commonly, their philosophy on licensing or OSS.
@@
4.2. RQ2: Processes for Addressing Licensing Problems
Here, we discuss how the study participants reported resolving licensing problems, including the involved roles, the adopted resolution process, and how participants understand licenses.
@@ 4.2.1. Roles
Legal professionals often participate in licensing compliance tasks, either in-house or as outside counsel (Wintersgill et al., 2024). We asked our participants whether they or their organization regularly worked with legal professionals or people otherwise trained in software licensing and/or copyright law. Eighteen participants responded “Yes,” 25 responded “No,” and 15 responded “I don’t know.” This indicates that expert-level licensing knowledge may not always be available to developers—or, in cases where it is, they may not be aware of it. This creates a particularly challenging situation since developers often do not have training associated with software licensing and find licensing tasks challenging (Almeida et al., 2019).
To understand developers’ perspectives on the issue in practice, we asked participants, “If an issue related to software licensing arises during software development, who is responsible for resolving that issue?” Participants could select developers, legal experts, an OSPO, or enter another role under “other.” As shown in Figure 6, respondents were divided on the issue. Although legal experts were most frequently cited as being responsible for addressing licensing issues (23), developers were also frequently listed (17). Notably, 5 respondents who answered that they did not regularly work with legal professionals and 4 respondents who reported not knowing if they or their organization regularly worked with legal professionals nonetheless selected legal experts as responsible for resolving licensing issues. Several respondents indicated they did not know whose responsibility it was (20), which may indicate a lack of experience or awareness in this topic. Developers rarely cited an OSPO, a group dedicated to managing activities related to open-source software, as responsible for addressing licensing issues (7). One respondent selected “other” and elaborated in a short answer that they were solely responsible for licensing in their case.
The number of respondents who viewed legal experts as responsible for addressing licensing issues (23) exceeds the number of developers who indicated that they regularly worked with such experts (18). This may indicate that the processes for addressing licensing issues are unstandardized and may not meet stakeholder expectations. We discuss the implications of this in Section 5.1.
Finding 6: Developers are divided on who is responsible for addressing licensing issues. Similarly sized groups chose legal experts and developers as responsible, while just as many did not know who was responsible.
@@ 4.2.2. Problem resolution process
To gauge the difficulty that developers had when addressing licensing issues, we asked survey participants to answer whether solutions to licensing issues were typically immediately apparent or if such issues required research. Participants were invited to elaborate. Of the 33 respondents who indicated that they had experienced licensing issues, only 6 respondents (10.34%) reported that solutions were readily apparent. Twenty-one respondents (36.21%) indicated that addressing licensing issues required additional research, while 6 respondents (10.34%) said that it depended on the nature of the problem.
@@
Many respondents then provided details on how they might address such problems. The most common such recourse was to involve a legal team in the process, as specified by 9 respondents (15.52%). Sometimes, this was represented as an interactive process involving both lawyers and developers: $R_{33}$ noted that “[t]he complexity depends on the type of license, usage rights, and legal implications, so it usually involves reviewing the specific terms of the license and possibly consulting legal or compliance experts.” In other cases, the process involved less “consulting” in favor of passing the issue to others, such as $R_{40}$, who said that issues went to “[m]y manager, prior [to contacting] the legal advisory team,” and $R_{38}$, who noted that “[f]urther investigation is passed to an expert in corporate law.” However, we note that such resources may not be available in all cases: for example, open-source development projects may have different resources available than corporate ones. Given that this was the most-reported strategy, this demonstrates a need for more accessible options for addressing licensing issues, or, at least, better adoption of such strategies. Four respondents (6.90%) stressed that an understanding of the involved software licenses and their terms or requirements is a necessary prerequisite to solving licensing problems. As $RI_{2}$ indicated in their survey response, to find solutions to licensing issues, “you need to fully understand the licensing in the first place.”
@@
Finding 7: When faced with licensing issues, developers’ most-reported strategy was escalating the issue to legal experts.
@@
The respondents also identified several specific strategies that could be deployed once a licensing issue was identified. Three respondents (5.17%) indicated that when issues with a dependency’s license arise, they might search for an alternative component with similar functionality. Two of these (3.45%) further noted that, as a last resort, they might write an internal alternative if a third-party solution were not feasible for licensing reasons. $R_{14}$ explains: “Usually we pivot to a different dependency with a compatible license or ask for license exceptions if possible. We had one case where we had to resort to doing an [in-house] development because the only library that matched the requirements was incompatible.”
@@
$RI_{6}$ noted that, concerning the timing of licensing tasks in the software development cycle, “the bulk happens at the first release…” This aligns with prior work showing that many organizations treat licensing as something to be addressed at the end of development, although such an approach introduces new challenges (Wintersgill et al., 2024).
@@
Some responses identified solutions that appeared only once (1.72%) in our results, including reverting to an older version of a dependency if that version had a more favorable license, purchasing a license when available, simply ceasing use of the affected component, or letting the issue go unaddressed. While the methods identified here align with previous work (Xu et al., 2023), our results show that no method was particularly favored, highlighting the need for licensing solutions and, by extension, tools that can offer solutions tailored to the specific circumstances of the software in question.
@@
Finding 8: Developers presented many strategies for addressing licensing issues at the dependency level, including changing dependencies, reverting to an old version, or implementing functionality themselves.

### v2

@@ 3.3. Interview design and analysis
To better understand the responses of the participants and gain further insight into their views on software licensing and related issues, we conducted semi-structured follow-up interviews over Zoom. At the end of the survey, we provided respondents with the option to leave an email address if they were willing to participate in a follow-up interview, which 21 participants utilized. We contacted all 21 of these individuals; of those, 7 participants responded and agreed to participate in follow-up interviews.
Each interview lasted 20 to 30 minutes and was recorded to facilitate analysis. The recordings were transcribed using Whisper (Whisper, ) and reviewed against the original recordings for accuracy by each researcher. Three reviewers assigned topic labels to the interviewee’s answers in each transcript, and all three reviewers then discussed the labels to reach a consensus.
We have assigned each interviewee an anonymous identifier to attribute the source of quotations (e.g., $RI_{4}$). As with the survey results, we have corrected grammar errors for readability and indicated deletions or additions by ellipses or brackets.
@@ 3.4. Participant demographics
Respondents came from 21 countries, with the largest representation from the United States (22, 38%), Germany (6, 10%), Bangladesh (5, 9%), and Colombia (5, 9%). Forty respondents (69%) reported that they have worked on proprietary software, 33 (57%) on open-source projects that relied on third-party components, and 14 (24%) on open-source projects with no dependencies. The presence of these different groups allows us to capture information on licensing tasks for different use cases. Most of the respondents described their primary role as programmer (32, 55%), with other common roles including researcher (10, 17%), project technical lead (5, 9%), tester (3, 5%), educator (3, 5%), and software architect (2, 3%). Development experience ranged from less than one year to 35 years, with an average of 9.9 years and a median of 7 years. Domains also varied broadly, ranging from contracted government work and healthcare to game development and banking. Almost two-thirds of the respondents had experience with web development (37, 64%). The respondents also developed libraries/frameworks (25, 43%), desktop applications (20, 34%), mobile applications (15, 26%), development tools (14, 24%), middleware (13, 22%), and AI-intensive systems (9, 16%). Although not extensive, the developers in the response pool represent a range of (self-reported) experiences and backgrounds, likely representing diverse experiences with software licensing.
We evaluated participants’ background with licensing by asking if they had any training or education in software licensing, and, if so, whether that training was formal (e.g., college classes, degree, or certification) or informal (self-learning, employer-provided, etc.). Thirty-five (60%) of our respondents reported that they had some kind of training. In total, 25 (43%) reported informal training, 4 (6.9%) reported formal training, 6 (10.3%) reported both formal and informal training, and 23 (39.7%) reported no licensing training. Despite low reported amounts of formal licensing training, almost all of these developers, as we will show in Section 4, have engaged in licensing tasks in some way.
As our previous work indicates, developers are on the front lines of license compliance (Wintersgill et al., 2024). This current work captures developers with a variety of backgrounds and experience levels to provide a realistic view at the current state of practice.
@@ 4. Results
In the following, we present the results of our study, which emerged from the survey responses and interviews.
4.1. RQ1: Developer Licensing Tasks
We discuss participants’ responses regarding the different software licensing tasks they perform and the challenges they face.
4.1.1. Most common tasks
Participants were presented with a list of software licensing-related tasks and asked to identify how often they participated in each task on a scale of “never,” “occasionally,” “frequently” or “all the time.” The tasks were derived from our previous experience working with licensing issues (such as license incompatibilities, license noncompliance, and problems with distribution) and from our prior work cataloging licensing bugs (Vendome et al., 2018). Notably, 55 respondents (95%) reported performing at least one licensing task at least occasionally. This suggests that most developers encounter licensing tasks at some point, further emphasizing the need to understand where the difficulties lie in these tasks. As shown in Figure 3, the tasks that developers reported performing the most frequently (i.e., frequently or all the time) were identifying the licenses of the software components they depend on, understanding the terms of use/service associated with the tools they use to develop software, and verifying that their software complies with all relevant licenses. Developers reported that they assigned a license to their own software slightly less often, though this was the task most reported by the respondents as being performed “all the time” (10). By contrast, most of the respondents indicated that they never spent time changing their software’s license (42) or acquiring or granting license exceptions (40), and slightly less than half reported that they never spent time identifying key license terms (28). While almost all developers reported performing licensing tasks at some point, all tasks were reported as being performed “never” or “occasionally” more than they were reported being performed “frequently” or “all the time.” This suggests that while most developers will need to address licensing tasks sometimes, developers perform licensing tasks relatively infrequently overall.
@@ 4.1.1. Most common tasks
Changing a software project’s license can be difficult (Wintersgill et al., 2024), particularly for open-source projects, but such changes occur (Vendome et al., 2017a). Few developers, however, reported having experience with changing the license of a software component, and no respondents reported performing this task “all the time.” Similarly, most developers reported that they had no experience in acquiring or granting license exceptions, in which an exception to the terms of a component’s license is granted to a particular licensee (for example, to grant additional permissions or remove restrictions that would otherwise be in place). License exceptions can be difficult to obtain from open source projects (Wintersgill et al., 2024) and may be a process with which developers are unfamiliar; indeed, this category received the most “I don’t know” responses, as shown in Table 1.
Almost half of the respondents reported that they never spent time identifying key terms from licenses. This may suggest that developers spend little time analyzing the contents of software licenses, or that they feel they understand available software licenses and how they interact, so they do not need to spend further time reviewing them.
@@
4.1.2. Most difficult tasks
Developers were also asked to rate the difficulty of licensing tasks using a 5-point Likert scale from “extremely difficult” to “extremely easy.” As shown in Figure 4, we observed that tasks such as identifying dependency licenses, understanding terms associated with development tools, and assigning a license were viewed as easiest, while changing a license, dealing with license exceptions, and identifying key license terms received more mixed responses. Notably, as Figure 4 shows, participants did not broadly accept any task as being easy: the results show mixed opinions at best for most of the tasks examined. These results confirm previous work showing that developers can find licensing tasks challenging (Almeida et al., 2017, 2019) while highlighting that some issues are viewed as being particularly difficult.
@@ 4.1.2. Most difficult tasks
Verifying whether software complies with all relevant licenses was cited as the most difficult task by respondents, being the task most frequently labeled “somewhat difficult” (25) as well as most frequently labeled “extremely difficult” (13). This is particularly notable, as compliance verification was reported as one of the most frequently performed tasks in the previous question. Identifying key license terms was also frequently cited as somewhat/extremely difficult. These trends in perceived difficulty are present across both developers who answered that they perform one or more licensing tasks “frequently” or “all the time” as well as developers who did not. Similarly, those who performed licensing tasks more frequently did not identify any tasks as significantly easier than those who did not perform licensing tasks frequently.
That developers struggle to understand licenses and their terms, and that verifying compliance proves a difficult task for them, is widely known in this area (Almeida et al., 2019; Gangadharan et al., 2008, 2012; Vendome et al., 2018). Our interviews provide additional context in this matter. For example, $RI_{1}$ details how different types of software can be more difficult to manage from a licensing perspective: “In order of [compliance] difficulty, you’d have the organizationally supported open source is the easiest, then proprietary software, then the non-organizationally supported open source because that one typically lacks consistent documentation, like consistent versioning, [etc. ].” $RI_{1}$ also contextualizes the difficulty in identifying licensing terms by explaining the practicality of it: “I find it difficult because of the length of licenses. If they’re not using a predefined license, like the Apache licenses or things like that […] it can be difficult, just because it’s several pages of licensing terms.” Beyond parsing many terms, it can also be necessary to interpret these terms as they apply to one’s specific circumstances. $RI_{7}$ identifies that “it’s very hard to make sure that your interpretation is the right one. And since we are talking in an international context, sometimes […] there are some cultural aspects that make you interpret the law in some ways. But when you are talking about something which is worldwide, it’s very difficult to interpret how the people that designed the rules thought about it.”
Finding 3: Verifying license compliance, while one of the most common licensing tasks, was also reported to be the most difficult. Identifying key license terms was also frequently perceived as difficult.
@@
Table 1. “I don’t know” / “I have no opinion” responses to survey questions about task frequency and difficulty
@@ 4.1.3. Licensing factors
To better understand the elements involved in the decision to license software, participants were asked to identify the factors they or their organizations consider when assigning licenses to software. Participants could select multiple such factors. As shown in Figure 5, developers most commonly cited legal considerations surrounding their software (30) as influencing the license assigned to their software. Together with the second most popular answer, organizational policies (24), this illustrates how licensing decisions are often made for non-technical reasons. The third most frequent answer, dependency licenses (23), shows how constraints from third-party components similarly affect such decisions. Only beyond this point do we observe specific considerations related directly to development-focused concerns, such as the software’s intended distribution (21) or intended use (19). Philosophical considerations, such as the views of the participant or their organization on ethical software distribution and use (13) or their views on the nature or purpose of OSS development (12), received fewer responses.
Overall, these results highlight a hierarchy of considerations when choosing a license: first, any legal obligations or organizational policies must be adhered to, then technical limitations based on the project and how it should be used, followed by philosophies regarding software and its reuse. It should be noted that, as participants were able to select multiple responses, a given respondent could highlight multiple of these groupings to be considered in tandem with each other, and for some respondents, the categories may overlap. However, the preeminence of non-engineering or non-usage requirements illustrates the first priorities respondents reported having when making licensing decisions. As such policies may vary among jurisdictions or among organizations, we suggest that licensing tools should accommodate organizational choices.
Finding 4: When considering how to license software, the surveyed developers appeared to first consider legal and organizational obligations before applying technical constraints based on the project and then, less commonly, their philosophy on licensing or OSS.
@@
4.2. RQ2: Licensing Issue Resolution
We discuss how study participants reported resolving licensing problems, including the involved roles, the adopted resolution process, and how the participants understand licenses.
@@ 4.2.1. Roles
Legal professionals often participate in licensing compliance tasks, either in house or as outside counsel (Wintersgill et al., 2024). We asked participants whether they or their organization regularly worked with legal professionals or people otherwise trained in software licensing and/or copyright law. Eighteen participants responded “Yes,” 25 responded “No,” and 15 responded “I don’t know.” This indicates that expert-level licensing knowledge may not always be available to developers (or they may not be aware of it). This creates a challenging situation, since developers often do not have training in software licensing and find licensing tasks challenging (Almeida et al., 2019).
To understand developers’ perspectives on the issue, we asked participants, “If an issue related to software licensing arises during software development, who is responsible for resolving that issue?” Participants could select developers, legal experts, OSPO, or enter another role under “other.” As shown in Figure 6, the respondents were divided on the issue. Although legal experts were listed most frequently (23), developers were also often chosen (17). Notably, 5 respondents who answered that they did not regularly work with legal professionals and 4 respondents who reported not knowing if they or their organization regularly worked with legal professionals nevertheless selected legal experts as responsible for resolving licensing issues, indicating that such answers were likely based on respondents’ beliefs regarding these responsibilities more than lived experience or known procedure. The participants’ answers may have depended on the resources available to them. Of the 40 respondents who indicated that they have worked on proprietary software, 20 said legal experts were responsible, while of the remaining 18 respondents who did not work on proprietary software, only 3 listed legal experts as responsible. Additionally, many respondents indicated that they did not know whose responsibility it was (20), which may indicate limited experience on this topic. Developers rarely cited an OSPO (Munir and Mols, 2021), a group dedicated to managing activities related to OSS, as responsible for addressing licensing issues (7). One respondent selected “other” and elaborated that they were solely responsible for licensing in their case.
The number of respondents who viewed legal experts as responsible for addressing licensing issues (23) exceeds the number of developers who indicated that they regularly worked with such experts (18). This may indicate that the processes for addressing licensing issues are not standardized and may not meet stakeholder expectations. We discuss the implications of this in Section 5.1.
@@
Finding 5: Developers are divided on who is responsible for addressing licensing issues. Similarly sized groups chose legal experts and developers as responsible, while just as many did not know who was responsible.
@@ 4.2.2. Problem resolution process
To gauge the difficulty that developers had when addressing licensing issues, we asked survey participants to answer whether solutions to licensing issues were typically immediately apparent or if such issues required research. The participants were invited to elaborate. Of the 33 respondents who indicated that they had experienced licensing issues, only 6 respondents (10.34%) reported that solutions were readily apparent. Twenty-one respondents (36.21%) indicated that addressing licensing issues required additional research, while 6 respondents (10.34%) said that it depended on the nature of the problem.
@@
Many respondents then provided details on how they might address such problems. The most common such recourse was to involve a legal team in the process, as specified by 9 respondents (15.52%). Sometimes, this was represented as an interactive process involving both lawyers and developers: $R_{33}$ noted that “[t]he complexity depends on the type of license, usage rights, and legal implications, so it usually involves reviewing the specific terms of the license and possibly consulting legal or compliance experts.” In other cases, the process involved less “consulting” in favor of passing the issue to others, such as $R_{40}$, who said that issues went to “[m]y manager, prior [to contacting] the legal advisory team,” and $R_{38}$, who noted that “[f]urther investigation is passed to an expert in corporate law.” However, as noted above, legal professionals may not always be available to developers. Given that this was the most-reported strategy, this demonstrates a need for more accessible options for addressing licensing issues, or, at least, better adoption of such strategies. Four respondents (6.90%) stressed that an understanding of the involved software licenses and their terms or requirements is a necessary prerequisite to solving licensing problems. As $RI_{2}$ indicated in their survey response, to find solutions to licensing issues, “you need to fully understand licensing in the first place.”
@@
Finding 6: When faced with licensing issues, developers’ most-reported strategy was escalating the issue to legal experts, but solutions often require additional research from developers and depend on the problem.
@@
The respondents also reported several specific strategies that could be deployed once a licensing issue was identified. Three respondents (5.17%) indicated that when issues with a dependency’s license arise, they might search for an alternative component with similar functionality. Two of these (3.45%) further noted that, as a last resort, they might write an internal alternative if a third-party solution was not feasible for licensing reasons. $R_{14}$ explained: “Usually we pivot to a different dependency with a compatible license or ask for license exceptions if possible. We had one case where we had to resort to doing an [in-house] development because the only library that matched the requirements was incompatible.”
@@
$RI_{6}$ noted that, concerning the timing of licensing tasks in the software development cycle, “the bulk happens at the first release.” This aligns with prior work showing that many organizations treat licensing as something to be addressed at the end of development, although such an approach introduces new challenges (Wintersgill et al., 2024).
@@
Solutions that appeared only once (1.72%) in our results included reverting to an older version of a dependency if that version had a more favorable license, purchasing a license when available, simply ceasing use of the affected component, or letting the issue go unaddressed. While the methods identified here align with previous work (Xu et al., 2023), our results show that no method was particularly favored, highlighting the need for licensing solutions and, by extension, tools that can offer solutions tailored to the specific circumstances of the software in question.
@@
Finding 7: Developers described disparate strategies for addressing licensing issues at the dependency level, including changing dependencies, reverting to an old version, and implementing functionality themselves, demonstrating a lack of uniform practice for handling licensing issues.